# Modeling extremal dependence in multivariate and spatial problems: a practical perspective

### v3

@@ 2.3 Parametric modeling of the extremal dependence
Theory. A non-exhaustive list of popular parametric models for the extremal dependence includes the asymmetric logistic (AL), pairwise beta (PB), tilted Dirichlet (TD), Hüsler-Reiss (HR), extremal-$t$ (ET) and extremal-skew-$t$ (EST) models (see e.g., Beranger and Padoan, 2015; Beranger et al., 2017).
The Poisson Point Process (PPP) is a useful characterization for statistical models used to represent the extremal dependence, when the latter is parametrized according to the angular measure and provided that it allows for a density $h(w\mid{\bm{\varphi}})$ on $\mathcal{S}_{d}$, where $\bm{\varphi}$} is a suitable vector of parameters (see e.g., Coles and Tawn, 1991; Engelke et al., 2015). In particular, let ${\bm{x}}_{1},\ldots,{\bm{x}}_{N}$ be i.i.d. observations from $F\in\mathcal{D}(G)$ and ${\bm{y}}_{1},\ldots,{\bm{y}}_{N}$ be their equivalent with unit Fréchet margins. The pseudopolar transformation $T$ is defined as $T({\bm{y}}_{i})=(r_{(i)},{\bm{w}}_{(i)})$ where $r^{(i)}=y_{i,1}+\cdots+y_{i,d}$ is the radial part and ${\bm{w}}_{(i)}={\bm{y}}_{i}/r_{(i)}$ is the angular part of ${\bm{y}}_{i}$, $i=1,\ldots,N$. The domain of attraction condition implies that if $r_{0}$ is a large threshold and $W_{r_{0}}=\{(r,{\bm{w}}):r>r_{0}\}$ is a set of points with a radial component larger than $r_{0}$, then the number of points falling in $W_{r_{0}}$ is described approximatively by a Poisson random variable with rate $1/\psi(W_{r_{0}})$, where $\psi=\xi\{T^{-1}(\cdot)\}$. Thus, conditionally on observing $k$ points $\{(r_{(i)},{\bm{w}}_{(i)}),i=1,\ldots,k\}$ with radial part greater than $r_{0}$, such points are approximately independent and generated from the density function $r^{-2}\mathrm{d}r\times H(\mathrm{d}{\bm{w}})/\psi(W_{r_{0}})$. The likelihood function of the exceedances can then be approximated as
$L({\bm{\varphi}}\mid(r_{(i)},{\bm{w}}_{(i)}),i=1,\ldots,k)$ $\approx\frac{e^{-\psi(W_{r_{0}})}(\psi(W_{r_{0}}))^{k}}{k!}\prod_{i=1}^{k}\frac{r_{(i)}^{-2}h({\bm{w}}_{(i)}\mid{\bm{\varphi}})}{\psi(W_{r_{0}})}$
@@
Software implementation. The routine fExtDep with argument method = "PPP" allows to estimate $\bm{\varphi}$} by likelihood maximisation. The model class for $h$ is specified with the argument model with the list of available options reported in Table 2. Some classes of models are characterised by an angular density $h({\bm{w}}\mid{\bm{\varphi}})$ that places a continuous mass only in the interior of $\mathcal{S}_{d}$ (PB, TD, HR), while others also allow to place densities and atoms on subspaces of $\mathcal{S}_{d}$. Beranger et al. (2017) describe an extended version of the ML method that allows for the estimation of these more complex models up to the three-dimensional case (AL, ET, EST) by exploiting the following approach.
@@
Application. We illustrate the use of fExtDep on the analysis of high levels of air pollution recorded in Leeds, UK, over the winter period November 1st to February 27/28th, between 1994 and 1998 (see Heffernan and Tawn, 2004). The pollution dataset contains the daily maximum of the five air pollutants: particulate matter (PM10), nitrogen oxide (NO), nitrogen dioxide (NO2), ozone (03), and sulfur dioxide (SO2), on the unit Fréchet scale (for a description of the transformation used, see Beranger and Padoan, 2015). Furthermore, the datasets PNS, NSN and PNN contains the angular components corresponding to the 100 largest radial components of the triplets of pollutants (PM10, NO, SO2), (NO2, SO2, NO) and (PM10, NO, NO2), respectively.
@@
initial value -118.269640iter 10 value -146.907539iter 10 value -146.907539final value -146.907539converged
R> f.et <- fExtDep(x = PNS, method = "PPP", model = "ET",+ par.start = c(0.5, 0.5, 0.5, 3), trace = 2 )
initial value -161.078869iter 10 value -225.654851iter 20 value -234.524086iter 30 value -235.285461iter 40 value -235.360084final value -235.382293convergedwhere the first call to the fExtDep routine fits the extremal-$t$ angular density (model = "ET"), to the pollution data (x = PSN), allowing for point masses at the corners by specifying the argument c, while the second call considers a density defined only in the interior of $\mathcal{S}_{d}$. The argument par.start gives starting values for the parameters and trace = 2 allows to monitor optimization progress. The optimization method is set by the argument optim.method ("BFGS" by default) on which basis the routine optim from the stats library is used. When method = "PPP" or "composite", fExtDep returns a list object of class ExtDep_Freq from which the estimates of the model parameters, their standard errors, the log-likelihood maximum and the Takeuchi Information Criterion can be extracted respectively with the est, StdErr, logLik and tic functions. The fitted model and estimation method can also be extracted using the model and method functions. fExtDep is first ran for the extremal-$t$ model using different values of $c$ ($0,0.05$ and $0.1$) to take into account point masses at the corners of $\mathcal{S}_{d}$. The TICs indicate that only assuming mass in the interior of $\mathcal{S}_{d}$ provides the best results (lowest TIC). A performance comparison between the PB, AL, TD, HR, ET and EST models defined on the interior only, is performed with TIC scores reported in Table 1 and as a result the extremal skew-$t$ model fits the data best (lowest TIC).
@@
Bayesian inference for the parameter $\bm{\varphi}$} is implemented through a model averaging approach (Sabourin et al., 2013) and available in the routine fExtDep when method = "BayesianPPP". Table 2 reports the implemented prior densities for each angular density model. In particular, $\phi(\cdot,a,b)$ stands for a one-dimensional Gaussian density with mean $a$ and variance $b$, $t(\cdot)$ is a transformation applied to the parameter (to map it to the real line). For each component of $\bm{\varphi}$}, the proposal density is also Gaussian, but of the form $\phi(t(\varphi_{j}^{\star});t(\varphi_{j}^{(i)}),\tt{MCpar})$, where $\varphi_{j}^{\star}$ and $\varphi_{j}^{(i)}$ are the proposed value and the value at the $i$th iteration of the algorithm of the $j$ element of $\bm{\varphi}$}, respectively, and MCpar is the variance term that needs to be specified by the user.
R> Hpar.hr <- list(mean.lambda = 0, sd.lambda = 3)R> PNS.hr <- fExtDep(x = PNS, method = "BayesianPPP", model = "HR",+ Nsim = 5e+4, Nbin = 3e+4, Hpar = Hpar.pb, MCpar = 0.35, seed = 14342)
|=====================================================================| 100%The above display illustrates how to perform Bayesian inference for the Hüssler–Reiss model (model = "HR") considering $50,000$ iterations (Nsim), a burn-in period of $30,000$ iterations (Nbin) while the hyper-parameters are set through Hpar and using the argument seed for reproducibility. Given that we have a Gaussian prior distribution for the model’s parameter on the log scale, setting the hyper-parameters mean.lambda=0 and sd.lambda=3 provides relatively non-informative prior information about the parameter. When method = "BayesianPPP", fExtDep returns an object of class ExtDep_Bayes from which the posterior mean, posterior standard deviation and Bayesian Information Criterion can respectively be extracted using the est, StdErr and bic functions. Again, the model and method function can be used to extract the model name and estimation procedure.
@@
The empirical mean (standard deviation) obtained from the approximate posterior distribution are $\hat{\lambda}_{1,2}=0.65(0.04),\hat{\lambda}_{1,3}=0.90(0.04)$ and $\hat{\lambda}_{2,3}=0.98(0.04)$, while the Bayesian Information Criterion (BIC) is 449.64-449.64. Visualization of the fitted angular density and data is obtained by calling the plot routine on the object of class ExtDep_Bayes with type = "angular" as default.
@@
R> # Function extracting the Generalized Pareto parameter estimates with aR> # threshold set at the 70% quantileR< est.fun <- function(x){+ x <- na.omit(x)+ unlist(evd::fpot(x, threshold = quantile(x, probs = 0.7))+ [c("threshold", "estimate")])+ }R> est <- apply(winterdat, 2, est.fun)R> # Transforming univariate data to unit Frechet scaleR> transform <- function (x, data, par){+ data <- na.omit(data)+ if(x > par[1]){+ emp.dist <- mean(data <= par[1])+ dist <- 1 - (1 - emp.dist) * max(0, 1 + par[3] * (x - par[1]) /+ par[2])^(-1 / par[3])+ }else{+ dist <- mean(data <= x)+ }+ return(-1/log(dist))+ }R> th <- c(95, 270, 110, NA, 95)R> Th <- sapply(c(1:3,5), function(x)+ transform(th[x], data = winterdat[, x], par = est[, x]))R> names(Th) <- colnames(winterdat[c(1:3, 5)])R> xl.PNS <- bquote("P(" * PM[10] ~ ">" ~ .(th[1]) * ", NO" ~ ">"+ ~ .(th[2]) * ", " * SO[2] ~ ">" ~ .(th[5]) * ")")R> # Compute and plot probabilities of joint exceedancesR> P.PNS <- pExtDep(q = Th[colnames(PNS)], type = "upper",+ method = "Parametric", model = "HR", par = PNS.hr\$stored.vals,+ xlab = xl.PNS)The left panel of Figure 3 provides the approximate posterior distribution of the tail probabilities in (2.10) for the extreme events. Note that a smoother output could be obtained by considering $50,000$ iterations (Nsim) and a burn-in period of $30,000$ iterations (Nburn). Formula (2.10) can also be used to derive a possible definition of the so-called joint return level, see Beranger and Padoan (2015) for details, i.e., the sequence of values $y_{j;p},j\in J\subset\{1,\ldots,d\}$, that satisfies the equation
@@ 2.4.1 Extremal dependence estimation via Bernstein polynomials
R> clust <- PAMmado\$clusteringR> d <- 5 # dimension = number of stationsR> stationsn <- matrix(NA, nclust, d)R> xx <- simplex(d = d, n = 15)R> fit <- list(length = nclust)R> est <- vector(length = nclust)R> set.seed(1)R> for(i in 1:nclust){+ # Randomly select 5 stations from cluster i+ stationsn[i, ] <- sample(which(clust == i), 5)+ data_tmp <- precip[, stationsn[i, ]]+ # Transform margins to unit Frechet scale+ data_uf_tmp <- trans2UFrechet(data_tmp, type = "Empirical")+ # Estimate Pickands dependence function+ fit[[i]] <- beed(data_uf_tmp, xx, d, "md", "emp", k = 7)+ # Extract extremal coefficient+ est[i] <- fit[[i]]\$extind+ }
@@
R> library(geosphere)R> pairs <- t(combn(92, 2))R> pairs.LLO <- cbind(pairs, lon[pairs[, 1]], lat[pairs[, 1]],+ lon[pairs[, 2]], lat[pairs[, 2]])R> pairs.dist <- apply(pairs.LLO, 1, function(x) distm(x = x[3:4],+ y = x[5:6], fun = distGeo)) / 1000R> pairs.EC <- pairs.ECBP <- vector(length = nrow(pairs))R> S <- simplex(d = 2, n = 49)R> for(i in 1:nrow(pairs)){+ # Transform margins to unit Frechet scale+ data.tmp <- trans2UFrechet(precip[,pairs[i, ]], type = "Empirical")+ # Estimate Pickands dependence function+ fit.tmp <- beed(data.tmp, S, 2, "md", "emp", k = 7)+ # Extract extremal coefficient+ pairs.ECBP[i] <- fit.tmp\$extind+ pairs.EC[i] <- 2 * madogram(S, data.tmp, "emp")[which(S[, 1] == 0.5)]+ }
@@ 2.4.2 Bernstein polynomials modeling and Bayesian nonparametric inference
where ${\bm{\eta}}_{\kappa}=(\eta_{0},\ldots,\eta_{\kappa-1})^{\top}$. Additionally, Marcon et al. (2016) established the conditions on the coefficients ${\bm{\eta}}_{\kappa}$ so that $H_{\kappa-1}$ is the distribution of a valid angular measure and proposed a Bayesian nonparametric procedure for the inference of both $A_{\kappa}$ and $H_{\kappa-1}$, which can then be used to compute an approximation of joint tail probabilities. The proposed method consists of three main steps: 1) the specification of a prior distribution for the polynomial order and coefficients $(\kappa,{\bm{\eta}}_{\kappa})$, 2) the derivation of the likelihood function, 3) the definition of a Markov Chain Monte Carlo (MCMC) algorithm for the posterior distribution computation. In particular, $\Pi(\kappa,{\bm{\eta}}_{\kappa})=\Pi({\bm{\eta}}_{\kappa}\mid\kappa)\,\Pi(\kappa)$, where $\Pi(\kappa)$ is a prior on the polynomial order (e.g., Poisson, negative Binomial, etc.) and $\Pi({\bm{\eta}}_{\kappa}\mid\kappa)=\Pi(\eta_{1},\ldots,\eta_{\kappa-2}\mid p_{1},p_{0},\kappa)\,\Pi(p_{1}\mid\kappa,p_{0})\,\Pi(p_{0})$ is a prior on the coefficients ${\bm{\eta}}_{\kappa}$, where $\Pi(p_{1}\mid\kappa,p_{0})$ and $\Pi(p_{0})$ are the priors on the coefficients representative of the atoms $\eta_{0}=p_{0}$ and $\eta_{\kappa-1}=1-p_{1}$ at the edges of the interval $[0,1]$. Such priors are specified as uniform distributions on suitable intervals, which have been chosen to ensure that the resulting Bernstein polynomial satisfies the constraint (C3). Specification of the prior $\Pi({\bm{\eta}}_{\kappa}\mid\kappa)$ induces also a prior on the coefficients ${\bm{\beta}}_{\kappa}$ of the corresponding polynomial $A_{\kappa}(t\mid{\bm{\beta}}_{\kappa})$, which automatically satisfy constraints (C1)–(C2). To deal with the fact at each MCMC iteration the dimension of ${\bm{\theta}}_{t}\kappa$ changes with $\kappa$, a trans-dimensional MCMC scheme is considered following Marcon et al. (2016). At the state $s$, $({\bm{\eta}}_{\kappa^{(s)}}^{(s)},\kappa^{(s)})$ is updated using the proposal distribution $q({\bm{\eta}}_{\kappa},\kappa\mid{\bm{\eta}}_{\kappa^{(s)}}^{(s)},\kappa^{(s)})=\Pi({\bm{\eta}}_{\kappa}\mid\kappa)q_{\kappa}(\kappa\mid\kappa^{(s)})$, where $q_{\kappa}(\kappa\mid\kappa^{(s)})$ is defined such that if $\kappa^{(s)}=3$, it places mass on $\kappa=4$ with probability $1$ and if $\kappa^{(s)}>3$ it places mass on $\kappa^{(s)}-1$ and $\kappa^{(s)}+1$ with equal probability. Finally, the likelihood function is defined as $\mathcal{L}({\bm{\vartheta}})=\prod_{i=1}^{m}\mathcal{L}({\bm{y}}_{i}\mid{\bm{\vartheta}})$, where ${\bm{\vartheta}}=({\bm{\theta}}_{1}^{\top},{\bm{\theta}}_{2}^{\top},{\bm{\beta}}_{\kappa}^{\top},\kappa)^{\top}$ and for any $y_{j}$ such that $(1+\gamma_{j}(y_{j}-\mu_{j})/\sigma_{j})>0$ with $j=1,2$,
@@
where $(Z_{1},Z_{2})$ are distributed as a bivariate GEV with unit-Fréchet margins, $z_{j}^{\star}=\sum_{i=1}^{M}\{1+\gamma_{i,j}(y_{j}^{\star}-\mu_{i,j})/\sigma_{i,j}\}^{1/\gamma_{i,j}}$ with $j=1,2$ and $\mathrm{B}(x\mid a,b)$, $x[0,1]$, denotes the distribution function of a beta random variable with shape parameters $a,b>0$ (see Marcon et al., 2016, for details).

### v5

@@ 2.3 Parametric modeling of the extremal dependence
Theory. The Poisson Point Process (PPP) is a useful characterization for statistical models used to represent the extremal dependence, when the latter is parametrized according to the angular measure and provided that it allows for a density $h(w\mid{\bm{\varphi}})$ on $\mathcal{S}_{d}$, where $\bm{\varphi}$} is a suitable vector of parameters (see, e.g., Coles and Tawn, 1991; Engelke et al., 2015). In particular, let ${\bm{x}}_{1},\ldots,{\bm{x}}_{N}$ be i.i.d. observations from $F\in\mathcal{D}(G)$ and ${\bm{y}}_{1},\ldots,{\bm{y}}_{N}$ be their equivalent with unit Fréchet margins. The pseudopolar transformation $T$ is defined as $T({\bm{y}}_{i})=(r_{(i)},{\bm{w}}_{(i)})$ where $r^{(i)}=y_{i,1}+\cdots+y_{i,d}$ is the radial part and ${\bm{w}}_{(i)}={\bm{y}}_{i}/r_{(i)}$ is the angular part of ${\bm{y}}_{i}$, $i=1,\ldots,N$. The domain of attraction condition implies that, if $r_{0}$ is a large threshold and $W_{r_{0}}=\{(r,{\bm{w}}):r>r_{0}\}$ is a set of points with a radial component larger than $r_{0}$, the number of points falling in $W_{r_{0}}$ is described approximatively by a Poisson random variable with rate $1/\psi(W_{r_{0}})$, where $\psi=\xi\{T^{-1}(\cdot)\}$. Thus, conditional on observing $k$ points $\{(r_{(i)},{\bm{w}}_{(i)}),i=1,\ldots,k\}$ with radial part greater than $r_{0}$, such points are approximately independent with the density function $r^{-2}\mathrm{d}r\times H(\mathrm{d}{\bm{w}})/\psi(W_{r_{0}})$. The likelihood function of the exceedances can then be approximated by
$L({\bm{\varphi}}\mid(r_{(i)},{\bm{w}}_{(i)}),i=1,\ldots,k)$ $\approx\frac{e^{-\psi(W_{r_{0}})}\{\psi(W_{r_{0}})\}^{k}}{k!}\prod_{i=1}^{k}\frac{r_{(i)}^{-2}h({\bm{w}}_{(i)}\mid{\bm{\varphi}})}{\psi(W_{r_{0}})}$
@@
Software implementation. A non-exhaustive list of popular parametric models for the extremal dependence includes the asymmetric logistic (AL), pairwise beta (PB), tilted Dirichlet (TD), Hüsler-Reiss (HR), extremal-$t$ (ET) and extremal-skew-$t$ (EST) models (see e.g., Beranger and Padoan, 2015; Beranger et al., 2017). The routine fExtDep with argument method = "PPP" allows to estimate $\bm{\varphi}$} by likelihood maximisation. The model class for $h$ is specified with the argument model with the list of available options reported in Table 2. Some classes of models are characterised by an angular density $h({\bm{w}}\mid{\bm{\varphi}})$ that places a continuous mass only in the interior of $\mathcal{S}_{d}$ (PB, TD, HR), while others also allow to place densities and atoms on subspaces of $\mathcal{S}_{d}$. Beranger et al. (2017) describe an extended version of a Maximum Likelihood (ML) method that allows for the estimation of these more complex models up to the three-dimensional case (AL, ET, EST) by exploiting the following approach.
@@
Application. We illustrate the use of fExtDep on the analysis of high levels of air pollution recorded in Leeds, UK, over the winter period (November 1st to February 28/29th) between 1994 and 1998 (see Heffernan and Tawn, 2004). The pollution dataset contains the daily maximum of the five air pollutants: particulate matter (PM10), nitrogen oxide (NO), nitrogen dioxide (NO2), ozone (03), and sulfur dioxide (SO2), on the unit Fréchet scale. For a description of the transformation used see Beranger and Padoan (2015). The datasets PNS, NSN and PNN contains the angular components corresponding to the 100 largest radial components of the triplets of pollutants (PM10, NO, SO2), (NO2, SO2, NO) and (PM10, NO, NO2), respectively.
@@
R> f.et <- fExtDep(x = PNS, method = "PPP", model = "ET",+ par.start = c(0.5, 0.5, 0.5, 3), trace = 2 )where the first call to the fExtDep routine fits the extremal-$t$ angular density (model = "ET"), to the pollution data (x = PSN), allowing for point masses at the corners by specifying the argument c, while the second call considers a density defined only in the interior of $\mathcal{S}_{d}$. The argument par.start gives starting values for the parameters and trace = 2 allows to monitor optimization progress. The optimization method is set by the argument optim.method ("BFGS" by default) on which basis the routine optim from the stats library is used. When method = "PPP" or "composite", fExtDep returns a list object of class ExtDep_Freq from which the estimates of the model parameters, their standard errors, the log-likelihood maximum and the Takeuchi Information Criterion (TIC) can be extracted respectively with the est, StdErr, logLik and tic functions. The fitted model and estimation method can also be extracted using the model and method S3 generics. fExtDep is first ran for the extremal-$t$ model using different values of $c$ ($0,0.05$ and $0.1$) to take into account point masses at the corners of $\mathcal{S}_{d}$. The TICs indicate that only assuming mass in the interior of $\mathcal{S}_{d}$ provides the best results (lowest TIC). A performance comparison between the PB, AL, TD, HR, ET and EST models defined on the interior only, is performed with TIC scores reported in Table 1 and as a result the extremal skew-$t$ model fits the data best (lowest TIC).
@@
Bayesian inference for the parameter $\bm{\varphi}$} is implemented through a model averaging approach (Sabourin et al., 2013) and available in the routine fExtDep using method = "BayesianPPP". Table 2 reports the implemented prior densities for each angular density model. Let $\phi(\cdot,a,b)$ denote a univariate Gaussian density with mean $a$ and variance $b$ and $t(\cdot)$ a transformation applied to the parameters to map them to the real line. For each component of $\bm{\varphi}$}, the proposal density is Gaussian of the form $\phi\{t(\varphi_{j}^{\star}\};t(\varphi_{j}^{(i)}),\tt{MCpar})$, where $\varphi_{j}^{\star}$ and $\varphi_{j}^{(i)}$ are the proposed value and the value at the $i$th iteration of the algorithm of the $j$ element of $\bm{\varphi}$}, respectively, and MCpar is the variance term that needs to be specified by the user.
R> Hpar.hr <- list(mean.lambda = 0, sd.lambda = 3)R> PNS.hr <- fExtDep(x = PNS, method = "BayesianPPP", model = "HR",+ Nsim = 3e3, Nbin = 1e3, Hpar = Hpar.hr, MCpar = 0.05, seed = 14342)The above display illustrates how to perform Bayesian inference for the Hüssler–Reiss model (model = "HR") with 3,000 iterations (Nsim), a burn-in period of 1,000 iterations (Nbin) while the hyper-parameters are set through Hpar and we set the random seed seed for reproducibility. Given that we have a Gaussian prior distribution for the model’s parameter on the log scale, setting the hyper-parameters mean.lambda=0 and sd.lambda=3 provides relatively non-informative prior information about the parameter. When method = "BayesianPPP", fExtDep returns an object of class ExtDep_Bayes from which the posterior mean, posterior standard deviation and Bayesian Information Criterion can respectively be extracted using the est, StdErr and bic functions. Note that here we have $2,000$ posterior draws corresponding to Nsim - Nbin. Again, the model and method function can be used to extract the model name and estimation procedure.
@@
The empirical mean (standard deviation) obtained from the approximate posterior distribution are $\hat{\lambda}_{1,2}=0.65(0.04),\hat{\lambda}_{1,3}=0.90(0.04)$ and $\hat{\lambda}_{2,3}=0.98(0.04)$, while the Bayesian Information Criterion (BIC) is 449.65-449.65. Visualization of the fitted angular density and data is obtained by calling the plot routine on the object of class ExtDep_Bayes with type = "angular" as default.
@@
R> # Function extracting the Generalized Pareto parameter estimates with aR> # threshold set at the 70% quantileR< est.fun <- function(x){+ x <- na.omit(x)+ unlist(evd::fpot(x, threshold = quantile(x, probs = 0.7))+ [c("threshold", "estimate")])+ }R> est <- apply(winterdat, 2, est.fun)R> # Transforming univariate data to unit Frechet scaleR> transform <- function (x, data, par){+ data <- na.omit(data)+ if(x > par[1]){+ emp.dist <- mean(data <= par[1])+ dist <- 1 - (1 - emp.dist) * max(0, 1 + par[3] * (x - par[1]) /+ par[2])^(-1 / par[3])+ }else{dist <- mean(data <= x)}+ return(-1/log(dist))+ }R> th <- c(95, 270, 110, NA, 95)R> Th <- sapply(c(1:3,5), function(x)+ transform(th[x], data = winterdat[, x], par = est[, x]))R> names(Th) <- colnames(winterdat[c(1:3, 5)])R> xl.PNS <- bquote("P(" * PM[10] ~ ">" ~ .(th[1]) * ", NO" ~ ">"+ ~ .(th[2]) * ", " * SO[2] ~ ">" ~ .(th[5]) * ")")R> # Compute and plot probabilities of joint exceedancesR> P.PNS <- pExtDep(q = Th[colnames(PNS)], type = "upper",+ method = "Parametric", model = "HR", par = PNS.hr\$stored.vals,+ xlab = xl.PNS)The left panel of Figure 3 provides the approximate posterior distribution of the tail probabilities in (2.10) for the extreme events. Note that a smoother output could be obtained by considering $3,000$ iterations (Nsim) and a burn-in period of $1,000$ iterations (Nburn). Formula (2.10) can also be used to derive a possible definition of the so-called joint return level, see Beranger and Padoan (2015) for details, i.e., the sequence of values $y_{j;p},j\in J\subset\{1,\ldots,d\}$, that satisfies the equation
@@ 2.4.1 Extremal dependence estimation via Bernstein polynomials
R> clust <- PAMmado\$clusteringR> d <- 5 # dimension = number of stationsR> stationsn <- matrix(NA, nclust, d)R> xx <- simplex(d = d, n = 15)R> fit <- list(length = nclust)R> est <- vector(length = nclust)R> set.seed(1)R> for(i in 1:nclust){+ # Randomly select 5 stations from cluster i+ stationsn[i, ] <- sample(which(clust == i), 5)+ data_tmp <- precip[, stationsn[i, ]]+ # Transform margins to unit Frechet scale+ data_uf_tmp <- trans2UFrechet(data_tmp, type = "Empirical")+ # Estimate Pickands dependence function+ fit[[i]] <- beed(data_uf_tmp, xx, d, "md", "emp", k = 7)+ # Extract extremal coefficient+ est[i] <- fit[[i]]\$extind}
@@
R> library(geosphere)R> pairs <- t(combn(92, 2))R> pairs.LLO <- cbind(pairs, lon[pairs[, 1]], lat[pairs[, 1]],+ lon[pairs[, 2]], lat[pairs[, 2]])R> pairs.dist <- apply(pairs.LLO, 1, function(x) distm(x = x[3:4],+ y = x[5:6], fun = distGeo)) / 1000R> pairs.EC <- pairs.ECBP <- vector(length = nrow(pairs))R> S <- simplex(d = 2, n = 49)R> for(i in 1:nrow(pairs)){+ # Transform margins to unit Frechet scale+ data.tmp <- trans2UFrechet(precip[,pairs[i, ]], type = "Empirical")+ # Estimate Pickands dependence function+ fit.tmp <- beed(data.tmp, S, 2, "md", "emp", k = 7)+ # Extract extremal coefficient+ pairs.ECBP[i] <- fit.tmp\$extind+ pairs.EC[i] <- 2 * madogram(S, data.tmp, "emp")[which(S[, 1] == 0.5)]}
@@ 2.4.2 Bernstein polynomials modeling and Bayesian nonparametric inference
where ${\bm{\eta}}_{\kappa}=(\eta_{0},\ldots,\eta_{\kappa-1})^{\top}$. Additionally, Marcon et al. (2016) established the conditions on the coefficients ${\bm{\eta}}_{\kappa}$ so that $H_{\kappa-1}$ is the distribution of a valid angular measure and proposed a Bayesian nonparametric procedure for the inference of both $A_{\kappa}$ and $H_{\kappa-1}$, which can then be used to compute an approximation of joint tail probabilities. The proposed method consists of three main steps: 1) the specification of a prior distribution for the polynomial order and coefficients $(\kappa,{\bm{\eta}}_{\kappa})$, 2) the derivation of the likelihood function, 3) the definition of a Markov Chain Monte Carlo (MCMC) algorithm for the posterior distribution computation. In particular, $\Pi(\kappa,{\bm{\eta}}_{\kappa})=\Pi({\bm{\eta}}_{\kappa}\mid\kappa)\,\Pi(\kappa)$, where $\Pi(\kappa)$ is a prior on the polynomial order (e.g., Poisson, negative Binomial, etc.) and $\Pi({\bm{\eta}}_{\kappa}\mid\kappa)=\Pi(\eta_{1},\ldots,\eta_{\kappa-2}\mid p_{1},p_{0},\kappa)\,\Pi(p_{1}\mid\kappa,p_{0})\,\Pi(p_{0})$ is a prior on the coefficients ${\bm{\eta}}_{\kappa}$, where $\Pi(p_{1}\mid\kappa,p_{0})$ and $\Pi(p_{0})$ are the priors on the coefficients representative of the atoms $\eta_{0}=p_{0}$ and $\eta_{\kappa-1}=1-p_{1}$ at the edges of the interval $[0,1]$. Such priors are specified as uniform distributions on suitable intervals, which have been chosen to ensure that the resulting Bernstein polynomial satisfies the constraint (C3). Specification of the prior $\Pi({\bm{\eta}}_{\kappa}\mid\kappa)$ induces also a prior on the coefficients ${\bm{\beta}}_{\kappa}$ of the corresponding polynomial $A_{\kappa}(t\mid{\bm{\beta}}_{\kappa})$, which automatically satisfy constraints (C1)–(C2). To deal with the fact at each MCMC iteration the dimension of ${\bm{\theta}}_{t}$ changes with $\kappa$, a trans-dimensional MCMC scheme is considered following Marcon et al. (2016). At iterate $s$, $({\bm{\eta}}_{\kappa^{(s)}}^{(s)},\kappa^{(s)})$ is updated using the proposal distribution $q({\bm{\eta}}_{\kappa},\kappa\mid{\bm{\eta}}_{\kappa^{(s)}}^{(s)},\kappa^{(s)})=\Pi({\bm{\eta}}_{\kappa}\mid\kappa)q_{\kappa}(\kappa\mid\kappa^{(s)})$, where $q_{\kappa}(\kappa\mid\kappa^{(s)})$ is defined such that if $\kappa^{(s)}=3$, it places mass on $\kappa=4$ with probability $1$ and if $\kappa^{(s)}>3$ it places mass on $\kappa^{(s)}-1$ and $\kappa^{(s)}+1$ with equal probability. Finally, the likelihood function is defined as $\mathcal{L}({\bm{\vartheta}})=\prod_{i=1}^{m}\mathcal{L}({\bm{y}}_{i}\mid{\bm{\vartheta}})$, where ${\bm{\vartheta}}=({\bm{\theta}}_{1}^{\top},{\bm{\theta}}_{2}^{\top},{\bm{\beta}}_{\kappa}^{\top},\kappa)^{\top}$ and for any $y_{j}$ such that $(1+\gamma_{j}(y_{j}-\mu_{j})/\sigma_{j})>0$ with $j=1,2$,
@@
where $(Z_{1},Z_{2})$ are distributed as a bivariate GEV with unit-Fréchet margins, $z_{j}^{\star}=\sum_{i=1}^{M}\{1+\gamma_{i,j}(y_{j}^{\star}-\mu_{i,j})/\sigma_{i,j}\}^{1/\gamma_{i,j}}$ with $j=1,2$ and $\mathrm{B}(x\mid a,b)$, $x\in[0,1]$, denotes the distribution function of a beta random variable with shape parameters $a,b>0$ (see Marcon et al., 2016, for details).